A rectangle with it dimensions in centimetres Length up(3y+2)cm and down(x+7)cm and breadth left 2xcm and right (x+y+1)cm (a) find the value of x and y (b) the dimension of the rectangle
Answer by greenestamps(13200)



Opposite sides of a rectangle have equal lengths, so

[1] 3y+2 = x+7
[2] 2x = x+y+1

To solve the problem....
(1) Solve [2] for x
(2) Substitute that value of x in [1] to find y
(3) Verify that your x and y satisfy the given conditions
(4) Answer the questions that are asked